# PyroGuardian: An IoT-Enabled System for Health and Location Monitoring in High-Risk Firefighting Environments

### v1

@@ 1. Introduction
Section two gives background regarding some wireless tools and sensors integrated into PyroGuardian. The following section describes a high-level overview of PyroGuardian’s model, workflow, and configuration. The “Implementation” section elaborates on PyroGuardian’s sensors and breakout boards. It will explain in detail the communication between each component. In addition, “Evaluation” presents our results, a comprehensive user study, and PyroGuardian’s performance based on our metrics, such as cost. Section six, Discussion, interprets our results and extends PyroGuardian to a broader scope. Finally, the conclusion states the future trends and ends the paper with our final remarks.
@@ 3.1. Data Model
Each PyroHelm and PyroStrap have a hardcoded firefighter ID for the PyroPortal to recognize the firefighter’s identity. Thus, ID pairing is necessary. For instance, a unit must wear the PyroHelm module with a firefighter ID 1 with the PyroStrap module that has the firefighter ID 1. In addition, they broadcast the firefighter ID along with sensor data with a long-range and low-power wireless protocol, LoRa, ranging 610 meters with our Adafruit boards.
We chose LoRa as some buildings are relatively large for Bluetooth or WiFi signals. Although the data rate of LoRa is minuscule, we can still continuously broadcast sensor data since we don’t stream pictures or audio. To explain PyroGuardian data model better, figure 2 shows its Chen Entity-Relationship model.
@@ 5.1. RQ1: PyroGuardian’s Cost
We believe the PyroHelm’s cost is at a feasible level. The PyroStrap contains fewer sensors compared to PyroHelmand costs less. We still assume a fixed price of \$10 for non-electronics materials. We listed PyroStrap’s equipment cost in table 2. The total cost of PyroStrap equipment is \$100.28.
@@
The total cost of our solution is \$497.21 with only one firefighter. Each additional firefighter will cost an extra \$254.78. Considering our solution with two firefighters costs approximately one thousand USD, we believe that PyroGuardian’s cost is reasonable.
We could not compare PyroGuardian’s cost to any other system in the literature as we did not find their cost sections. We then looked for the pricing for similar systems in the industry. We found the Microsoft-supported 3AM Innovation’s Florian system that provided features such as firefighter location, geofencing, and 3D Satellite Maps (flo, 2024). We did not see Florian providing health monitoring services like PyroGuardian. They operate on an annual subscription basis, and only the subscription for the commander costs \$1000 (flo, 2024). For each individual, it costs \$240 yearly, and each vehicle \$500 yearly (flo, 2024). We believe PyroGuardian provides comparable pricing to Florian, such as our one-time fee of \$254.78 per firefighter instead of \$240 yearly, and a one-time fee of \$242.4 for only commander instead of \$1000.
@@ 5.2. RQ2: Requirement Satisfaction
The 34 participant firefighters watched the PyroPortal dashboard from a large-screen TV and the burning building from outside. After the two units had completed the demo, we asked the rest of the spectator firefighters to complete the UMUX-lite survey and describe their experience with the incident. The UMUX-lite survey has two questions, but we added four side questions that ask the participants to explain their reasoning with the two main questions. The two main questions are on a linear scale of 1 to 5, one representing strongly disagree, three neutral, and five strongly agree. We obtained 28 surveys in total.
The first main question we asked was if PyroGuardian’s capabilities meet their requirements. The answers we received to our survey are in table 3. We received an average score of 4.5. We consider this score satisfactory to show that PyroGuardian covers firefighters’ use case.
Our side questions are in table 3. For the first and last side questions, we received recommendations to enable backup and logging for accountability and compliance for PyroPortal. PyroPortal still contains valuable information, such as the firefighter list and the incident addresses. We can integrate a cloud database and add a SIM card to the tablet to enable an internet connection. We also plan on incorporating a logging solution to keep track of each firefighter’s actions in future iterations. This method can provide accurate records to ICs to improve accountability.
A feature that participants stated was missing was the thermal imaging streaming. Our thermal imaging is currently non-functional and a placeholder. They indicated that ICs should see the thermal video stream from firefighters’ gas masks since the smoke causes the environment to be pitch-black. LoRa is not capable of video streaming. However, future technologies, such as WiFi HaLow, have higher bandwidth that may support audio or video streaming in a 1-kilometer range (Knyazev et al., 2017). This architecture change would reduce the range from 10 km to 1 km as HaLow has less range than LoRa (Knyazev et al., 2017). However, we believe 1 km would suffice for most first responder indicators.
@@ 5.3. Misc RQ2: PyroGuardian’s Range
We focused on the range although other metrics were available, such as PyroStrap and PyroHelm’s battery lives. We calculated the battery life to be around 8 hours, which we believe would suffice for most incidents.
Some disaster zones are relatively large. The maximum transmission range between the IC and the firefighter will influence the total length a unit can move before losing connection. Thus, we maximized the range between the tablet and firefighter’s wearable modules with LoRa. We measured the maximum range of the LoRa Adafruit RFM95 Transceiver at 915 MHz with objects and walls between us to be 610 meters. We believe that distance is an acceptable range for most residential buildings to allow the IC to stay safe while the firefighters enter the dangerous area. We can further increase this distance with a LoRa mesh network in future iterations.
@@ 5.4. RQ3: User Convenience
The second main question we asked in table 3 was if PyroGuardian was easy to use. The average score we received was 4.18. Although this score is lower than the first main question, we believe it is acceptable. We can improve the user experience in future iterations by integrating tooltips and tutorials.
@@ 6. Related Work
Work suggesting wearable sensors to monitor firefighter health metrics already exists in the literature and industry (Meina et al., 2020). Meina et al. suggest using heart rate variability (HRV) and accelerometric data to monitor cardiovascular responses to situations (Meina et al., 2020). Their system combines HRV and an accelerometer to measure stress levels during extreme conditions, but it does not deduce environmental information, such as a firefighter’s ambient overheating. PyroPortal calculates derived values, such as stress level and toxicity.
@@
Outdoor navigation is included in our system with GPS, as the commander can accurately track units and see the direction they’re facing. This information is helpful for navigating a unit toward a particular destination. Amanatiadas et al. tackled indoor tracking and developed a system specifically for first responders (Amanatiadis et al., 2011). Their system uses wearable modules, which contain a foot-mounted inertial measurement unit (IMU), a digital camera, and a radio frequency (RF) identification device for first responders (Amanatiadis et al., 2011). IMU sensor utilizes the accelerometer data originating from walking patterns to perform dead-reckoning operations (Amanatiadis et al., 2011). It fuses with the camera’s 3D positioning algorithm and the location estimate from the Wi-Fi device (Amanatiadis et al., 2011). This solution focuses on Simultaneous localization and mapping (SLAM) but has a lower range with Wi-Fi.
@@
A work that fills the gap of vital information monitoring and still uses localization comes from O’Flynn et al. as they developed a system named SAFESENS (O’Flynn et al., 2018). It aims to provide an efficient rescue operation and first responder location tracking (O’Flynn et al., 2018). SAFESENS has eight building blocks, including UWB localization access points, an occupancy detection camera, a vital signs monitoring system, an explosive gas detector, and a firefighter tracking node (O’Flynn et al., 2018). The firefighter carries a smartphone that harvests data and sends it to a server for processing (O’Flynn et al., 2018). One of the modules is responsible for localization and uses stationary UWB anchor points (O’Flynn et al., 2018). The vital signs monitoring system performs advanced tasks, such as detecting blood composition changes (O’Flynn et al., 2018). Their system requires previously configured UWB anchor points for localization, which may not be possible in disaster scenarios. It does not have GPS as well to track firefighters outdoors.
@@
Instead of performing first responder localization, Fruhling et al. developed the system REaCH to solely include real-time health monitoring through first responders’ wearable modules (Fruhling et al., 2020). This system is similar to ours as they also use a dashboard to help the incident commander evaluate if his units need to leave the scene due to health complications (Fruhling et al., 2020). However, their system does not include indoor or outdoor location tracking or geofencing.
@@
Besides the work we found on Google Scholars, there are also companies in this area. 3M partnered with Microsoft and developed a similar platform (mic, 2021). The platform, Florian Enroute, provides ICs a 3-D street view and vehicle tracking capabilities, which calculates their ETAs (flo, [n.d.]). It allows ICs to use voice commands to navigate their platform and assign vehicles to specific locations (flo, [n.d.]). Although their product has unique features, such as voice commands or geo-fencing, they do not cover vital health monitoring. Their specialization is more on location tracking, while we aim to unify these features in PyroGuardian. However, their voice command capability may enable ICs to perform mission control operations while driving safely and inspire us to include it in our future work.
@@ 7. Discussion
This section discusses PyroGuardian’s possible application areas and analyzes PyroGuardian’s features regarding fire squads’ safety.
7.1. Analysis of Firefighter Safety with PyroGuardian
The fire service community recognizes cardiac issues as common health risks in the field (Fahy et al., 2020). There are cardiovascular risks for firefighters as their duties have strenuous physical demands, sympathetic nervous system activation, increased thermal burden, and environmental exposure to smoke pollutants (Lefferts et al., 2021). These factors also make cancer a well-recognized risk for retired firefighters, as the Association of Fire Fighters stated there were more than 130 firefighter cancer deaths in 2019 (Fahy et al., 2020). During their operations, they experience these factors, requiring a professional to monitor their vitals. Therefore, our PyroStrap’s pulse, heart rate, and oximeter sensors help the IC keep track of his units’ vitals. If the commander realizes the vitals are in the danger range, they can signal the PyroHelm module to command the unit to leave the scene for rest or medical care. Although PyroGuardian does not entirely remove the injury risk, the system would mitigate it during operations and reduce cancer rates among retired units.
@@ 7.2. Extended Use Cases of PyroGuardian
Falls in the elderly are a significant source of injury that can result in a disability or hospitalization (Pfortmueller et al., 2014). Serious injuries, such as fractures and traumatic brain injuries, occur in approximately 10% of falls among people older than 65 (Pfortmueller et al., 2014). With the fall detection feature, PyroHelm can help elder care facilities. The fall detection wearable can be in the pocket of an elder. The module will not predict or stop falls or slips from happening. However, it can alert a manager if the elder is incapacitated to notify emergency personnel.
@@ 7.4. Future Work
Although GPS fixes are rare in indoor environments, PyroGuardian does not provide indoor localization. Some techniques fuse GPS and IMU information to predict location (Beauregard and Haas, 2006). One such method is pedestrian dead reckoning (PDR) (Beauregard and Haas, 2006). It starts with a known position, and the next position displacements are added from the acceleration data (Beauregard and Haas, 2006). The IMU can calculate the following positions from the acceleration data, but the error will accumulate. Until the error drift becomes a significant value, the system can operate with IMU. Once the next GPS signal is received, the system can correct itself.

### v2

@@ 1. Introduction
Section two gives background regarding some wireless tools and sensors integrated into PyroGuardian. The following section describes a high-level overview of PyroGuardian’s model, workflow, and configuration. The “Implementation” section elaborates on PyroGuardian’s sensors and breakout boards. It will explain in detail the communication between each component. In addition, “Evaluation” presents our results, a comprehensive user study, and PyroGuardian’s performance based on our metrics, such as cost. Then, we describe the related papers to our work. Section seven, Discussion, interprets our results and extends PyroGuardian to a broader scope. Finally, the conclusion states the future trends and ends the paper with our final remarks.
@@ 3.1. Data Model
Each PyroHelm and PyroStrap has a hardcoded firefighter ID for the PyroPortal to recognize the firefighter’s identity. Thus, ID pairing is necessary. For instance, a unit must wear the PyroHelm module with a firefighter ID 1 with the PyroStrap module that has the firefighter ID 1. In addition, all wearables broadcast their firefighter ID along with sensor data with a long-range and low-power wireless protocol, LoRa, ranging 610 meters with our Adafruit boards.
We chose LoRa as some buildings are relatively large for Bluetooth or WiFi signals. Although LoRa’s data rate is minuscule, it can still fit our sensor data. PyroGuardian doesn’t stream pictures or audio, as LoRa isn’t the appropriate tool for it. To explain PyroGuardian data model better, figure 2 shows its Chen Entity-Relationship model.
@@ 3.2. Data Processing at PyroPortal
3.3. Analysis of Firefighter Safety with PyroGuardian
The fire service community recognizes cardiac issues as common health risks in the field (Fahy et al., 2020). There are cardiovascular risks for firefighters as their duties have strenuous physical demands, sympathetic nervous system activation, increased thermal burden, and environmental exposure to smoke pollutants (Lefferts et al., 2021). These factors also make cancer a well-recognized risk for retired firefighters, as the Association of Fire Fighters stated there were more than 130 firefighter cancer deaths in 2019 (Fahy et al., 2020). During their operations, they experience these factors, requiring a professional to monitor their vitals. Therefore, our PyroStrap’s pulse, heart rate, and oximeter sensors help the IC keep track of their units’ vitals. If the commander realizes the vitals are in the danger range, they can signal the PyroHelm module to command the unit to leave the scene for rest or medical care. Although PyroGuardian does not entirely remove the injury risk, the system would mitigate it during operations and reduce cancer rates among retired units.
@@ 5.1. RQ1: PyroGuardian’s Cost
The PyroStrap contains fewer sensors compared to PyroHelm and costs less. We still assume a fixed price of \$10 for non-electronics materials. We listed PyroStrap’s equipment cost in table 2. The total cost of PyroStrap equipment is \$100.28.
@@
The total cost of our solution is \$497.21 with only one firefighter. Each additional firefighter will cost an extra \$254.78.
@@ 5.2. RQ2: Requirement Satisfaction
The 34 participant firefighters watched the PyroPortal dashboard from a large-screen TV and the burning building from outside. After the two units had completed the demo, we asked the spectator firefighters to complete the UMUX-lite survey.
The UMUX-lite survey has two questions and is on a linear scale of 1 to 5, one representing strongly disagree, three neutral, and five strongly agree. Scoring and interpreting a UMUX-lite survey is similar to a System Usability Scale (SUS) (Vandereecken, 2021). Researchers normalize the average of the two survey questions’ scores to get a final result between 0 and 100, where higher scores indicate better-perceived usability (Vandereecken, 2021). After normalization, we obtained 86.8/100.
We also added the four custom side questions in table 4 to our UMUX-lite survey that asked the participants to explain the reasoning behind their ratings. We obtained 28 surveys in total. Finally, we interviewed each firefighter individually to learn more about their perspectives and recommendations.
The first question from our UMUX-lite survey was if PyroGuardian’s capabilities meet their requirements. The answers we received to our survey are in table 3. We received an average score of 4.5.
For the first and last side questions, we received recommendations to enable backup and logging for accountability and compliance for PyroPortal. PyroPortal still contains valuable information, such as the firefighter list and the incident addresses. We can integrate a cloud database and add a SIM card to the tablet to enable an internet connection. We also plan on incorporating a logging solution to keep track of each firefighter’s actions in future iterations. This method can provide accurate records to ICs to improve accountability.
In the interviews, a feature that participants stated was missing was the thermal imaging streaming. Our thermal imaging is currently non-functional and a placeholder. They indicated that ICs should see the thermal video stream from firefighters’ gas masks since the smoke causes the environment to be pitch-black. LoRa is not capable of video streaming. However, future technologies, such as WiFi HaLow, have higher bandwidth that may support audio or video streaming in a 1-kilometer range (Knyazev et al., 2017). This architecture change would reduce the theoretical range from 10 km to 1 km as HaLow has less range than LoRa (Knyazev et al., 2017). However, we believe 1 km would suffice.
@@ 5.3. Misc RQ2: PyroGuardian’s Range
We focused on the range although other metrics were available, such as PyroStrap and PyroHelm’s battery lives. We calculated the battery life to be around eight hours for each wearable.
Some disaster zones are relatively large. The maximum transmission range between the IC and the firefighter will influence the total length a unit can move before losing connection. Thus, we maximized the range between the tablet and firefighter’s wearable modules with LoRa. We measured the maximum range of the LoRa Adafruit RFM95 Transceiver at 915 MHz with objects and walls between us to be 610 meters.
@@ 5.4. RQ3: User Convenience
The second question in our UMUX-lite survey in table 3 was if PyroGuardian was easy to use. The average score we received was 4.18.
@@ 6. Related Work
Work suggesting wearable sensors to monitor firefighter health metrics already exists in the literature and industry (Meina et al., 2020). However, these projects are on a much smaller scale, and none of them offer a unified dashboard that covers health, environment, and location monitoring without prior configuration. Meina et al. suggest using heart rate variability (HRV) and accelerometric data to monitor cardiovascular responses to situations (Meina et al., 2020). Their system combines HRV and an accelerometer to measure stress levels during extreme conditions, but it does not deduce environmental information, such as a firefighter’s ambient overheating. PyroPortal calculates derived values, such as stress level and toxicity.
@@
Outdoor navigation is included in our system with GPS, as the commander can accurately track units and see the direction they’re facing. This information is helpful for navigating a unit toward a particular destination. Amanatiadas et al. tackled indoor tracking and developed a system specifically for first responders (Amanatiadis et al., 2011). Their system uses wearable modules, which contain a foot-mounted inertial measurement unit (IMU), a digital camera, and a radio frequency (RF) identification device for first responders (Amanatiadis et al., 2011). IMU sensor utilizes the accelerometer data originating from walking patterns to perform dead-reckoning operations (Amanatiadis et al., 2011). It fuses with the camera’s 3D positioning algorithm and the location estimate from the Wi-Fi device (Amanatiadis et al., 2011). This solution focuses on Simultaneous localization and mapping (SLAM), has a lower range with Wi-Fi, and doesn’t cover health monitoring.
@@
A work that fills the gap of vital information monitoring and still uses localization comes from O’Flynn et al. as they developed a system named SAFESENS (O’Flynn et al., 2018). It aims to provide an efficient rescue operation and first responder location tracking (O’Flynn et al., 2018). SAFESENS has eight building blocks, including UWB localization access points, an occupancy detection camera, a vital signs monitoring system, an explosive gas detector, and a firefighter tracking node (O’Flynn et al., 2018). The firefighter carries a smartphone that harvests data and sends it to a server for processing (O’Flynn et al., 2018). One of the modules is responsible for localization and uses stationary UWB anchor points (O’Flynn et al., 2018). The vital signs monitoring system performs advanced tasks, such as detecting blood composition changes (O’Flynn et al., 2018). Their system requires previously configured UWB anchor points for localization, which may not be possible in disaster scenarios. It also does not monitor environmental variables or have GPS to track firefighters outdoors.
@@
Instead of performing first responder localization, Fruhling et al. developed the system REaCH to solely include real-time health monitoring through first responders’ wearable modules (Fruhling et al., 2020). This system is similar to ours as they also use a dashboard to help the incident commander evaluate if their units need to leave the scene due to health complications (Fruhling et al., 2020). However, their system does not include indoor or outdoor location tracking or geofencing.
@@
Besides the work we found in the literature, there are also companies in this area. 3M partnered with Microsoft and developed a similar platform (mic, 2021). The platform, Florian Enroute, provides ICs a 3-D street view and vehicle tracking capabilities, which calculates their ETAs (flo, [n.d.]). It allows ICs to use voice commands to navigate their platform and assign vehicles to specific locations (flo, [n.d.]). Although their product has unique features, such as voice commands or geo-fencing, they do not cover vital health or environmental monitoring. Their specialization is more on location tracking, while we aim to unify these features in PyroGuardian. However, their voice command capability may enable ICs to perform mission control operations while driving safely and inspire us to include it in our future work.
@@ 7. Discussion
This section discusses PyroGuardian’s evaluation results, explains its further use cases, provides techniques for PyroGuardian’s data security, and analyzes the future work items.
7.1. Interpretation of Results
Considering our solution with two firefighters costs approximately one thousand USD, we believe that PyroGuardian’s cost is reasonable. However, we could not compare PyroGuardian’s cost to any other system in the literature as we did not find their cost sections. We then looked for the pricing for similar systems in the industry. We found the Microsoft-supported 3AM Innovation’s Florian system that provided features such as firefighter location, geofencing, and 3D Satellite Maps (flo, 2024). We did not see Florian providing health monitoring services like PyroGuardian. They operate on an annual subscription basis, and only the subscription for the commander costs \$1000 (flo, 2024). For each individual, it costs \$240 yearly, and each vehicle \$500 yearly (flo, 2024). We believe PyroGuardian provides comparable pricing to Florian, such as our one-time fee of \$254.78 per firefighter instead of \$240 yearly, and a one-time fee of \$242.4 for only commander instead of \$1000.
Besides PyroGuardian’s cost, we believe our 4.5 score from our UMUX-lite survey’s first question is satisfactory to show that PyroGuardian covers firefighters’ use case. Our 4.18 score from our UMUX-lite survey’s second question is lower than the first question. We can improve the user experience in future iterations by integrating tooltips and tutorials.
Overall, we received 86.8/100 from our UMUX-lite survey, and its benchmarks follow the SUS, which has an average score of 68 (Sauro, [n.d.]). A SUS score between 84.1 and 100 is the best imaginable result and is in the 96-100 percentile (Sauro, [n.d.]). The Net Promoter Score (NPS) is a metric used to measure customer satisfaction and loyalty by asking customers how likely they are to recommend a product, and our result converts to their highest score too: Promoter (Sauro, [n.d.]). Promoters are most likely to recommend the product/website/app to a friend (Sauro, [n.d.]).
Besides customer satisfaction metrics, we also think PyroGuardian’s eight hours battery life and distance of 610 meters distance are acceptable for most incidents and residential buildings to allow the IC to stay safe away from the disaster zone. We can further increase this distance with a LoRa mesh network in future iterations.
@@ 7.2. Extended Use Cases of PyroGuardian
Falls in older people are a significant source of injury that can result in a disability or hospitalization (Pfortmueller et al., 2014). Serious injuries, such as fractures and traumatic brain injuries, occur in approximately 10% of falls among people older than 65 (Pfortmueller et al., 2014). With the fall detection feature, PyroHelm can help elder care facilities. The fall detection wearable can be in the pocket of an elder. The module will not predict or stop falls or slips from happening. However, it can alert a manager if the elder is incapacitated to notify emergency personnel.
@@ 7.4. Future Work
GPS fixes are rare in indoor environments, so PyroGuardian does not provide indoor localization. Some techniques fuse GPS and IMU information to predict location (Beauregard and Haas, 2006). One such method is pedestrian dead reckoning (PDR) (Beauregard and Haas, 2006). It starts with a known position, and the next position displacements are added from the acceleration data (Beauregard and Haas, 2006). The IMU can calculate the following positions from the acceleration data, but the error will accumulate. Until the error drift becomes a significant value, the system can operate with IMU. Once the next GPS signal is received, the system can correct itself.